### v2

# Tightly Coupled SLAM with Imprecise Architectural Plans

Abstract
Robots navigating indoor environments often have access to architectural plans, which can serve as prior knowledge to enhance their localization and mapping capabilities. While some SLAM algorithms leverage these plans for global localization in real-world environments, they typically overlook a critical challenge: the “as-planned” architectural designs frequently deviate from the “as-built” real-world environments. To address this gap, we present a novel algorithm that tightly couples LIDAR-based simultaneous localization and mapping with architectural plans under the presence of deviations. Our method utilizes a multi-layered semantic representation to not only localize the robot, but also to estimate global alignment and structural deviations between “as-planned” and “as-built” environments in real-time. To validate our approach, we performed experiments in simulated and real datasets demonstrating robustness to structural deviations up to cm and . On average, our method achieves less localization error than baselines in simulated environments, while in real environments, the “as-built” 3D maps show lower average alignment error.
Paper Video: https://www.youtube.com/watch?v=eEZPkcpjWlM
I Introduction
Prior information from architectural plans can be used to enhance the localization and mapping accuracy of mobile robots. However, real-world buildings rarely match their plans perfectly due to construction tolerances and modifications. Incorporating such imprecise prior information into robot navigation pipelines can introduce systematic errors, potentially damaging the localization and mapping accuracy, instead of helping to improve it. To address this issue, it is critical to tightly couple the Simultaneous Localization and Mapping (SLAM) pipeline with the architectural plans. And for this, one needs a unified representation for both “as-planned” and “as-built” environments, that captures not only the geometry but also the architectural semantics. This is what will allow us to match structural elements between the “as-planned” and “as-built” instances and estimate the deviations between the two.
Although there are various environment representation techniques such as occupancy grids [1], surface maps [2], octomap [3] etc, none of them explicitly models the semantic and hierarchical information of environment, which is needed to identify the deviated structural elements. Recent approaches such as 3D scene graphs [4, 5] or Situational Graphs (S-Graphs) [6, 7], represent a robot’s environment in a compact and hierarchical manner, encoding high-level semantic abstractions (for example, walls and rooms) and their relationships (e.g., a set of walls forms a room). Herein, S-Graphs extend 3D scene graphs by merging geometric models of the environment generated by SLAM approaches with 3D scene graphs into a multi-layered jointly optimizable factor graph. This representation, combined with the prior information extracted from architectural plans, can be used to provide fast and efficient localization.
Informed S-Graphs (iS-Graphs) [8] further extend S-Graphs by using architectural plans to provide accurate localization over the resulting hierarchical factor graphs. However, its success is based on the assumption that there are no deviations between the “as-built” and the “as-planned”. In reality, this is never the case, and the building elements exhibit certain deviations with respect to their planned geometries.
The main contribution of this paper is a novel method capable of coupling architectural plans (“as-planned”) and SLAM (“as-built”) data even in the presence of deviations, as shown in Fig. 1. We call this algorithm Deviations Informed Situational Graphs or diS-Graphs in short.
In summary, diS-Graphs performs three interconnected tasks simultaneously:
Coupling of SLAM factor graph with architectural plans to match structural elements (walls, rooms) between planned and built environments.
Globally localizing the robot in imprecise architectural plans.
Detecting and estimating the deviations between the matched structural elements of “as-planned” and “as-built” environments in real-time.
II Related Works
Most localization techniques using prior information from architectural plans assume that the environments are built precisely according to the plans. One of the most commonly used localization techniques in 2D metric prior maps is Monte Carlo Localization (MCL) [9], [10] but it is not scalable to large-scale complex environments. Boniardi et al [11] use a technique that scales to more complex environments by aligning a scan-based map with CAD-based floor plans. OGM2PGM [12] also scales to larger environments by converting the 2D floor plan to an occupancy grid map (OGM) and using a pose-graph map (PGM) to localize the robot. UKFL [13] further enhances the localization accuracy using an unscented Kalman filter to localize the robot in 3D metric meshes. Recent techniques such as [14] exploit neural networks to localize the robot using an implicit neural representation of the floor plans. All of the above mentioned techniques primarily rely on geometric information, not using any possible semantic information available in the architectural plans, limiting their ability to reason about the environment beyond geometric features. This geometric-only approach fails to leverage crucial semantic cues like room types, door locations, and functional spaces that could help disambiguate similar-looking areas and improve localization accuracy. In addition, inaccuracies or outdated information in the floor plan can significantly affect the performance of these methods.
To address these limitations, semantic-based localization techniques, such as Mendez et al. [15] use semantic cues from architectural plans and sensor information to improve localization accuracy. Boniardi et al. [16] exploit the semantics of the room in architectural plans to do robot localization by matching the detected rooms from sensor data. Wang et al. [17] leverage prelabeled architectural features, such as wall intersections and corners, as landmarks in floor plans, and match them with detection from sensor data to jointly perform mapping and localization. Zimmerman et al. [18], [19] use high-level semantic information in floor plans, derived from object detection, along with geometric data from 2D LiDAR to perform long-term robot localization in floor plans. Huan et al. [20] convert architectural plans into semantically enriched point cloud maps, followed by a coarse-to-fine localization process using ICP. Gao et al. [21] used neural networks to detect vertical elements from floor plans to do LiDAR based localization. These methods are prone to inaccuracies due to misidentification and errors in the pose estimate of semantic elements. Moreover, they treat semantic elements in isolation, failing to leverage the rich contextual information embedded in their spatial and functional relationships.
Recent work such as Shaheer et al. [8] exploit the topological relationship between semantic elements to localize the robot with respect to architectural plans. However, all the above mentioned approaches assume no deviations between the architectural plans and the actual environment.
Some recent works leverage imprecise floor plans for localization. Boniardi et al. [22] integrate localization techniques to take advantage of the information embedded in the CAD drawing, and the real-world observations acquired during navigation, which may not be reflected in the floor plan. Li et al. [23] presented a 2D LiDAR-based localization system in imprecise floor plans using stochastic gradient descent (SGD) with a scan matching algorithm. Chan et al. [24] presented a 2D LiDAR-based localization in floor plans that integrates SLAM with MCL. Blum et al. [25] use neural networks for feature segmentation and combine them with LiDAR data for localization in imprecise floor plans. While these works demonstrate robot localization in inaccurate floor plans, they lack the capability to identify element-wise deviations between ”as-planned” and ”as-built” environments.
Despite advances in robot localization using imprecise floor plans, no existing approach combines localization with the estimation of structural deviations between “as-planned” and “as-built” environments, limiting the ability to assess construction accuracy in real time. In this work, we address this limitation by presenting a unified framework that enables both global localization and deviation estimation.
III System Architecture
We propose an algorithm that tightly couples SLAM and architectural plans and jointly optimizes them.
III-A Factor Graph Modelling
Our algorithm models both SLAM and architectural plans as hierarchical factor graphs (see Fig. 3) composed of the following elements:
Architectural Graph (A-Graph). Three-layered hierarchical factor graph model of the geometry, semantics, and topology of an environment, generated from its architectural plan. It models the environment “as-planned” by the architect.
Situational Graph (S-Graph). Four-layered hierarchical optimizable factor graph built online from 3D LiDAR and odometry measurements [6], [7] which models the “as-built” environment. It also includes the keyframes in addition to the geometry, semantics, and topology of the environment.
Deviations Informed Situational Graph (diS-Graph). diS-Graph is the result of couppling both A-Graph and S-Graph.
These graphs are organized in the following layers:
Origins Layer. Both S-Graph and A-Graph are defined in their own reference frame. The reference frame of S-Graph and A-Graph are called and respectively. Both the origin frames are coupled via a Transformation factor called .
Keyframes Layer. Only present in S-Graphs, this layer contains as nodes the robot poses .
Walls Layer. In an A-Graph, this layer’s nodes encode two semantic entities, namely wall-surfaces and walls . Every wall-surface is defined by four coefficients representing its normal orientation and distance to origin [6]. We assume that each wall has two planar wall-surfaces with opposite orientations and the separation between them is equal to the width of the wall. In S-Graphs, this layer contains only the wall-surfaces extracted from 3D LiDAR scans. The keyframes that observe such wall-surfaces are linked to them through pose-plane constraints. This layer of both graphs is coupled via wall deviation factors in diS-Graph.
Rooms Layer. In an A-Graph, this layer also encodes two semantic entities, namely Rooms consisting of four wall-surfaces and Doorways . Two rooms constrain a doorway, and a room is constrained by four walls. In S-Graphs, this layer contains rooms comprising either four wall-surfaces or two wall-surfaces, and does not contain doorways. This layer of both graphs is coupled via room deviation factors in diS-Graph.
Floors Layer. In both A-Graph and S-Graph this layer consists of a floor center node represented as , constraining all rooms present at that particular floor level.
More details on the type of constraints between the different elements of the graphs can be found in [8].
III-B The Algorithm
Our algorithm has two stages, as shown in Fig. 2. In Stage-A, the A-Graph is first matched and coupled with the S-Graph, and then jointly optimized. Stage-A runs until it finds the initial match between the two graphs, and provides the initial estimates of both the transformation between graphs and potential deviations between their elements. Afterwards, in Stage-B the joint graph is continuously refined through incremental matching and optimization as the robot navigates the environment, incorporating newly detected semantic entities. This tight coupling enables continuous refinement of both the robot’s localization and the detected structural deviations during navigation.
Both stages of our algorithm consist of multiple processes, described briefly below and detailed in their respective sections:
Graph Matching (Section IV). The Global Graph Matching in Stage-A provides the first unique match, when it exists, between the S-Graph and A-Graph at room and wall-surface levels, accounting for potential deviations. In Stage-B, the Local Graph Matching extends the previously matched elements in the diS-Graph with newly detected elements following an incremental approach.
Graph Coupling (Section. V). It is performed for the candidates matched by graph matching. In Stage-A, graph coupling registers the origins of two graphs along with the semantic coupling (wall-surfaces and rooms), and explicit mapping of the deviation factors. Stage-B only performs the semantic coupling that incorporates the newly observed entities with explicit deviation factors.
Graph Optimization (Section. VI). It optimizes the coupled graph. Stage-A involves two steps: alternating optimization and joint optimization. Stage-B only does joint optimization. Alternating optimization estimates two types of variables: 1. The ransformation between the origins of the S-Graph and the A-Graph yielding global localization, 2. The possible deviations between the matched graph entities. In joint optimization the coupled graphs are jointly optimized given the initial global transformation and deviation estimates.
IV Graph Matching
Our graph matching extends the method presented in [8]. In [8], a top-down potential candidate search between the A-Graph and S-Graph is performed by leveraging their hierarchical structures. To assess the overall consistency of each generated candidate, two verification steps are applied iteratively. First, the consistency of the node type and graph structure is verified. Second, geometric consistency (i.e. norm) is maintained over a certain consistency threshold [26].
It is worth noting that symmetries may occur due to a lack of information as the robot has not visited the whole environment and thus the algorithm requires more information to provide a unique match, or when a large part of a building is symmetric and a unique match could never be found.
IV-A Deviation-based Candidates Search
The presence of deviations generates geometric inconsistencies that affect the aforementioned candidates’ checks. Concretely, a deviation in the position of a wall-surface implies a slight deviation in the center of its parent room as well. To handle this, we propose a two-stage search algorithm where we first search for non-deviated wall-surfaces followed by the inclusion of those that are deviated.
Non-Deviated Elements Search To handle potential deviations, our method relaxes the matching criteria by decreasing the consistency thresholds at each level. First, we apply relaxed consistency thresholds for the generation of room-to-room and room-to-wall-surface match candidates, to account for the induced room-center inconsistencies. Then, to exclude deviated wall-surfaces, we increase the threshold for wall-surface-to-wall-surface candidates.
Deviated Elements Detection To identify the deviated wall-surfaces which were not matched in the first stage but are connected to already matched rooms, we decrease the consistency threshold at wall-surface-to-wall-surface level.
To further speed up the candidate search, we incorporate the following information: Orphan Wall-Surfaces: We utilize wall-surfaces in the S-Graph without a parent room for the assessment of the geometrical consistency of the final match candidates at the wall-surface level. Ground Orientation: We exploit the ground plane normal in the A-Graph and the S-Graph, only allowing candidates with z-axis rotations.
Finally, the geometric consistency score provides a quantification of the probability of the deviation for each room and wall-surface, which is further used in the Graph Coupling step (Section V).
IV-B Incremental Matching
To enhance the efficiency of the Graph Matching algorithm in [8], we propose an incremental approach with two stages (associated with stages A and B of the systems architecture of Fig. 2), namely Global Graph Matching (Fig. 4 left) and Local Graph Matching (Fig. 4 right), each executing the two previously described deviation-based candidate search stages. Until a first unique match has been found, the Global Graph Matching is executed for every new observation in the S-Graph. Afterward, the Local Graph Matching is executed every time the diS-Graph is updated with newly observed rooms and wall surfaces. Here, already-matched elements are excluded from candidate generation, and each assessment of intra-level consistency considers the previously matched elements at the corresponding level.
V Graph Coupling
Origin Coupling. We couple the origins of two graphs by introducing a transformation factor . The cost function is defined as:
Here and are the origins of the A-Graph and S-Graph respectively, and is the transformation between them. stands for the covariance of the cost, and it is always assigned a high value to estimate the transformation factor accurately.
Semantic Coupling. We couple the wall-surfaces and rooms of the A-Graph and the S-Graph to estimate deviations between the two graphs by introducing deviation factors as follows:
Room Coupling: To estimate the deviation in the pose of a matched room between the two graphs, we define a deviation factor between the two rooms as , where the cost function is defined as:
Here and are rooms of A-Graph and S-Graph, each consisting of a set of four wall-surfaces. is the covariance associated with the cost function depending on the probability of deviation estimated by the matching of the graph (Section IV). Rooms with a higher probability of deviation assigned by graph matching have a higher covariance assigned to their cost function than rooms with lower deviation probability. If all matched rooms have the same deviation probability, they are assigned lower uniform covariances.
Wall-Surface Coupling: After coupling the rooms of the A-Graph and the S-Graph, we then couple the wall-surfaces of the coupled rooms. We define the deviation factor between two wall surfaces as . The cost function to estimate the deviation value is defined as:
Here , where and are the normal orientation and distance of a plane in the A-Graph. Similarly, is a plane of the S-Graph with respect to the S-Graph origin. is the covariance associated with the cost function. Like rooms, wall-surfaces with a higher deviation are assigned higher covariances, and the ones with a lower deviation are assigned lower uniform covariances.
VI Graph Optimization
Before the S-Graph and A-Graph are coupled, the S-Graph state at time can be defined as:
Similarly, the A-Graph state is defined as:
where are the robot poses at selected keyframes in the S-Graph frame of reference, are the plane coefficients of the and wall-surfaces of the S-Graph and A-Graph respectively, contains the poses of the and four-wall rooms of the S-Graph and A-Graph respectively. are the poses of the two-wall rooms in the S-Graph. contains the poses of doorways of the A-Graph, are the floors levels, and models the drift between the odometry frame and the S-Graph reference frame . If at time there is no match obtained between the A-Graph and S-Graph we perform single S-Graph optimization as detailed in [7].
Alternating Optimization. Alternating optimization is further performed in two steps as follows:
Transformation Estimation: Upon receiving the match (Section IV) and performing graph coupling (Section V), we combine Eq. 4 and Eq. 5 to make a global state with additional transformation factor . represents the transformation between the origins of the A-Graph and the S-Graph. It is important to note that at this stage, the deviated wall-surface and room are not included for estimating .
Deviation Estimation: After optimizing we already have an initial guess of the transformation between the A-Graph and the S-Graph, and we can incorporate the deviated wall-surface and room entities into the graph with appropriate deviation factors. Our state then becomes where are the deviation factors between wall-surfaces and are the deviation factors between rooms. When optimizing we keep constant to obtain a good initial estimation of the deviation between the matched deviated entities.
Joint Optimization. Finally, after getting the initial estimates of the transformation between the origins and the deviations between the semantic entities, we optimize the whole state to simultaneously estimate the position of each semantic entity, deviations, and the transformation between the two graphs.
VII Experimental Evaluation
VII-A Methodology
Setup. We evaluated our algorithm in both simulated and real environments. Both simulated and real experiments were performed using a laptop computer with an Intel i9-11950H (8 cores, 2.6 GHz) with 32 GB of RAM.
Baselines. We have selected the following LiDAR-based baselines for comparisons due to their suitability, their reported results, and the availability of their code: AMCL [9], UKFL [13], OGM2PGM [12], IR-MCL [14], and iS-Graphs [8]. Although some works do localization in imprecise architectural plans, the absence of open-source implementations prevents direct comparison with these methods. As each selected baseline takes a different map input for localization, we used 2D occupancy grid maps for AMCL and OGM2PGM, 3D meshes for UKFL and IR-MCL, and A-Graphs for iS-Graphs and our method diS-Graphs, respectively.
Simulated Datasets. We validate the algorithms in five simulated datasets named SE1 to SE5. To record the datasets, we use the Gazebo physics simulator to recreate the robot, its sensors (LiDAR), and the 3D indoor environments obtained from actual architectural plans. We report the absolute trajectory error (ATE) compared with the available ground truth trajectory. We also report the localization convergence success rate of all methods.
Real Datasets. We collected data with a legged robot equipped with a Velodyne VLP-16 LiDAR, at five different construction sites (RE1 to RE5), with existing architectural plans. In real experiments, we report the Root Mean Square Error (RMSE) between estimated 3D maps and ground truth architectural plans due to the absence of pose ground truth. The RMSE, while a mapping metric, primarily demonstrates the localization accuracy of all baselines, as mapping accuracy inherently reflects pose estimation quality. Furthermore, we report all methods’ convergence rates and convergence and computation times.
Deviations. Note that, given the absence of actual ground truth deviations in a real environment with respect to the plans, we explicitly deviate the wall-surfaces entities in the architectural plans to have a ground truth estimate of the deviations for both simulated and real datasets. For all datasets, we performed five separate tests introducing uniform random wall-surface deviations in architectural plans ranging from to in translation, and and in rotations. Moreover, to test the robustness of our algorithm against deviations, we conducted 25 experiments per dataset with varying deviations.
Ablation. We conducted ablation studies on two key components of our algorithm. 1. Covariance assignment: As mentioned in the graph matching and graph merging (Section. IV & V), we assign covariances to the deviation factors of the semantic elements, based on their deviation likelihood. Here, we analyze the effect of assigning equal covariances, referred to as uniform covariances (UC), to both deviated and non-deviated elements. 2. Alternating optimization: In graph optimization (Section. VI) we discuss the use of alternating optimization for simultaneous localization and deviation estimation. Here, we study the performance when using only a single optimization cycle (SO).
VII-B Results and Discussion
Absolute Trajectory Error. Table I shows the average ATE for all baselines and our diS-Graphs in the presence of deviations between “as-planned” and “as-built” environments. Our method shows an error reduction of around compared to AMCL, compared to OGM2PGM, compared to IR-MCL, compared to UKFL, and a reduction of over iS-Graphs.
Bold values are the best and the second best are underlined. ‘-’ refers to an unsuccessful run.
| Dataset | ||||||
| Method | ATE [cm] | |||||
| SE1 | SE2 | SE3 | SE4 | SE5 | Avg. | |
| AMCL [9] | 17.2 | 20.1 | 22.4 | 19.9 | ||
| UKFL [13] | 12.6 | 15.3 | 8.7 | 11.1 | 9.1 | |
| OGM2PGM [12] | 15.2 | 18.1 | 10.7 | 10.3 | 14.3 | 13.7 |
| IR-MCL [14] | 14.7 | 6.4 | 9.6 | 28.4 | 18.8 | 15.5 |
| iS-Graphs [8] | 5.4 | 6.7 | 16.6 | 4.6 | 9.5 | 8.5 |
| diS-Graphs (UC) | 6.2 | 7.3 | 13.7 | 8.4 | 8.6 | 8.8 |
| diS-Graphs (SO) | 4.4 | 6.1 | 15.2 | 7.7 | 6.2 | 7.9 |
| diS-Graphs (Ours) | 3.3 | 4.1 | 6.4 | 4.4 | 5.7 | 4.8 |
Fig. 6(a) summarizes the ATE performance of all the baseline algorithms. AMCL shows the highest median ATE and a relatively narrow distribution, indicating consistently high error rates. UKFL and OGM2PGM demonstrate moderate performance with similar median ATEs, although OGM2PGM shows a wider range of errors. IR-MCL exhibits the largest variability, suggesting inconsistent performance in different scenarios. AMCL, OGM2PGM, IR-MCL, UKFL and iS-Graphs assume no deviations between “as-planned” and “as-built” environments. These methods attempt to match laser scans with an incorrect reference map, leading to consistent misalignment. Therefore, they show higher ATEs across datasets, indicating their vulnerability to architectural deviations. diS-Graphs can detect and estimate the deviations between “as-planned” and “as-built” environments, resulting in lower ATE than other baselines.
Point Cloud Alignment Error. Table II shows the RMSE of the point clouds with respect to the ground truth for all methods and ours. In case of deviations in construction from the plans, diS-Graphs shows better accuracy than AMCL, better than OGM2PGM, better than IR-MCL, and better than iS-Graphs. Although UKFL’s average error is equal to diS-Graphs’, it has a very low convergence rate (see Table III) rendering the comparison unfair. Moreover, it cannot estimate the deviations between “as-planned” and “as-built” environments. Fig. 6(b) summarizes the performance of all algorithms in real environments. AMCL exhibits the highest median and widest interquartile range, indicating greater variability in performance. IR-MCL presents a large spread of results, while OGM2PGM shows moderate performance with a smaller range of variability compared to AMCL and IR-MCL. Although UKFL and diS-Graphs show the lowest median RMSE, suggesting superior accuracy, the comparison with UKFL is not fair because of its low convergence rate. Because of our simultaneous estimation of deviations and initial transformation, we can not only simultaneously globally localize the robot but also estimate the deviations between semantic elements of “as-planned” and “as-built” environments moreover improving the overall accuracy compared to other algorithms.
| Dataset | ||||||
| Method | Point Cloud RMSE [cm] | |||||
| RE1 | RE2 | RE3 | RE4 | RE5 | Avg | |
| AMCL[9] | 0.48 | 0.88 | 0.31 | 0.56 | ||
| UKFL [13] | 0.31 | 0.20 | 0.26* | |||
| OGM2PGM[12] | 0.46 | 0.57 | 0.33 | 0.49 | 0.55 | 0.48 |
| IR-MCL[14] | 0.51 | 0.76 | 0.36 | 0.54 | ||
| iS-Graphs [8] | 0.27 | 0.29 | 0.23 | 0.33 | 0.28 | |
| diS-Graphs (UC) | 0.27 | 0.32 | 0.23 | 0.34 | 0.29 | |
| diS-Graphs (SO) | 0.27 | 0.30 | 0.23 | 0.33 | 0.28 | |
| diS-Graphs (ours) | 0.25 | 0.28 | 0.20 | 0.31 | 0.26 | |
* Omitted due to low convergence.
Convergence Rate. Our approach demonstrates superior convergence rates across both simulated and real environments (Table III). In simulations, our method achieves a convergence rate better than AMCL, better than UKFL, and better than OGM2PGM. While our convergence rate matches that of IR-MCL in simulations, IR-MCL exhibits inconsistent performance in terms of ATE and requires retraining for each new dataset. Similarly, in real environments (RE1-RE4), our method maintains the highest convergence rate among all baselines, leveraging its ability to detect deviated elements in the environment. The only exception is in RE5, where our algorithm fails due to insufficient room detection for effective graph matching, because of noisy sensor data. Moreover, AMCL and UKFL’s heavy reliance on accurate initial position estimates leads to degraded convergence rates in environments with complex geometries. OGM2PGM’s accuracy suffers particularly in symmetric environments. Our method overcomes these limitations through its robust deviation detection capabilities, though it requires a sufficient number of distinguishable rooms to function effectively.
| Dataset | ||||||||||
| Convergence Rate [%] | ||||||||||
| Method | SE1 | SE2 | SE3 | SE4 | SE5 | RE1 | RE2 | RE3 | RE4 | RE5 |
| AMCL [9] | 0 | 100 | 0 | 80 | 20 | 80 | 100 | 100 | 0 | 0 |
| UKFL [13] | 80 | 80 | 0 | 40 | 80 | 60 | 0 | 60 | 0 | 0 |
| OGM2PGM [12] | 80 | 100 | 100 | 80 | 80 | 60 | 80 | 100 | 60 | 80 |
| IR-MCL [14] | 100 | 100 | 100 | 20 | 100 | 60 | 100 | 20 | 0 | 0 |
| iS-Graphs [8] | 60 | 20 | 100 | 80 | 100 | 60 | 20 | 40 | 60 | 0 |
| diS-Graphs | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 0 |
| Dataset | ||||||||||
| Convergence Time [s] | Computation Time [ms] | |||||||||
| Method | RE1 | RE2 | RE3 | RE4 | RE5 | RE1 | RE2 | RE3 | RE4 | RE5 |
| AMCL [9] | 24 | 89 | 29 | 2 | 2 | 2 | ||||
| UKFL [13] | 126 | 8 | 104 | 119 | ||||||
| OGM2PGM [12] | 16 | 38 | 27 | 35 | 79 | 2 | 2 | 2 | 2 | 2 |
| IR-MCL [14] | 16 | 26 | 13 | 92 | 90 | 89 | ||||
| iS-Graphs [8] | 155 | 101 | 78 | 139 | 57 | 78 | 78 | 64 | ||
| diS-Graphs | 81 | 43 | 46 | 139 | 56 | 77 | 76 | 70 | ||
| Seq. Len. [s] | 657 | 170 | 488 | 657 | 559 | 657 | 170 | 488 | 657 | 559 |
Convergence and Computation Time. Table IV shows the convergence time for each algorithm. The convergence time is the time it takes the algorithm to globally localize. IR-MCL and OGM2PGM have the best convergence time. iS-Graphs and diS-Graphs have considerably longer convergence times because they need to detect a certain number of rooms in the environment for graph matching to find a unique match. In addition, the previous version of graph matching struggled to resolve symmetries during the matching process, resulting in longer convergence times. However, the modifications in the graph matching algorithm (Section IV) proposed in this work improve the symmetry resolution ability and reduce the convergence time by almost .
Table IV shows the computation time for each algorithm. The computation time is the time used for each pose update. On average, the computation time of our algorithm is milliseconds, showing its real-time performance. Our computation time is the best compared to other 3D algorithms. OGM2PGM and AMCL have the best computation time because, unlike others, they process 2D information.
Deviation Estimation. Fig. 8 shows the amount of deviation our algorithm can correctly estimate in real environments. The minimum detectable deviation is bounded by our LiDAR sensor’s accuracy. With our LiDAR having an accuracy of [27] and considering that of measurements fall within three standard deviations () of the mean, our system can reliably detect deviations only when they exceed . The maximum translational and rotational deviation in wall-surfaces our algorithm can detect accurately is and respectively. Fig. 5 shows several qualitative results from our diS-Graphs method for real experiments. The robot can successfully localize, map, and estimate deviations in these environments. Note that we only show the rooms and walls for better understanding, and all the other semantic elements and the robot are not shown. We do not show RE5 in Fig. 5, as the robot could not localize itself in this sequence.
Robustness against Deviations. Figure 7 shows the effect of deviations on the localization performance of iS-Graphs and diS-Graphs. The results consistently show that diS-Graphs, which incorporates deviation detection and modeling, maintains a relatively stable ATE even as the deviation increases from cm to cm. In contrast, iS-Graphs, which lacks explicit deviation modeling capabilities, exhibits a clear upward trend in ATE as deviations increase.
Ablation Study. Table I shows that associating ‘uniform covariance’ in simulated datasets (diS-Graphs (UC)) the algorithm cannot differentiate between deviated and non-deviated elements which results in poor pose estimation. Fig. 7 demonstrates that iS-Graphs with uniform covariance shows higher error variability compared to our method which uses deviation probability-based covariance assignment. In some cases, the use of uniform covariances results in even worse performance than iS-Graphs. Similarly, when using single optimization (diS-Graphs (SO)) instead of alternating optimization, the algorithm cannot differentiate between the transformation between two graphs and the deviations between their elements, resulting in higher ATE as shown in Table I. Table II shows the ablation of uniform covariances and single optimization in real-world datasets. Using uniform covariances (diS-Graphs (UC)) and single optimization (diS-Graphs (SO)) results in lower mapping accuracy due to the algorithm’s inability to accurately map the deviated elements and differentiate between initial transformation and deviations simultaneously.
VIII Conclusions and Future Work
This paper presents a novel approach to tightly couple SLAM with imprecise architectural plans. Our algorithm establishes direct correspondences between structural elements in “as-built” environments and their “as-planned” counterparts, enabling simultaneous robot localization and deviation estimation between the two representations. By detecting and estimating the deviations between “as-built” and “as-planned” environments through tight coupling, our method outperforms the current best approach with 43% better localization in simulations, 7% improved mapping accuracy in real environments, and enhanced robustness to architectural deviations. Additionally, our algorithm provides an estimate of existing deviations up to in translation and in rotation. Our algorithm is limited by the need to have enough distinctive semantic elements (i.e. wall-surfaces and rooms) to provide a unique match. As future work, we plan to add the ability to detect and match more semantic elements, which will translate into an improvement in the convergence rate and deviation detection range. Moreover, to gain flexibility, we plan to improve the matching process by adding the ability to detect rooms consisting of more than four walls.